OpenStudy (anonymous):

Medal Solve the following system of equations by any method. x + y = 32 x + z = 39 y + z= 47

OpenStudy (lynfran):

which method u prefer to use

OpenStudy (anonymous):

addition

OpenStudy (lynfran):

what i meant was elimination , substitution ... matrix

OpenStudy (anonymous):

substitution

hero (hero):

Subtract the 2nd from the 1st to get y-z=-7

hero (hero):

Then you'll have : y-z =-7 y+z=47 Add the equations together to eliminate z then solve for y.

hero (hero):

That's the easiest approach here.

OpenStudy (anonymous):

y = 0 z = 0

OpenStudy (anonymous):

-7 + 47 = 40

hero (hero):

You should have gotten 2y=40 not y=0

OpenStudy (anonymous):

oh sorry i subtracted

hero (hero):

Yes I know. Be careful

OpenStudy (anonymous):

whats the next step

OpenStudy (anonymous):

@LynFran do you understand this

OpenStudy (lynfran):

here ur substitution method and yes i know all 3 method thats y i ask which u prefer ok we can rewrite the 3rd equation as y=47-z we can also rrewrite the 2nd equation to get x=39-z now subs the 2 equations about into the 1st equation to get x+y=32 39-z+47-z=32 86-2z=32 86-32=2z 54=2z 27=z now sub z=27 into the 3rd equation above to get y+z=47 y+27=47 y=47-27 y=20 now sub y=20 into 1st equation to get x+y=32 x+20=32 x=32-20 x=12

hero (hero):

Elimination is the simplest and most efficient method to Use here.

OpenStudy (lynfran):

yea i agree but @harz360 prefer substitution

hero (hero):

The goal is to solve using the least amount of steps

OpenStudy (anonymous):

when i see all the steps thats when i understand it

hero (hero):

It's a matter of efficiency not preference.

OpenStudy (texaschic101):

I would have taken the first 2 equations...and eliminated the x's...then go from there

OpenStudy (lynfran):

me too

hero (hero):

Exactly .

OpenStudy (anonymous):

but thanks a lott @LynFran

OpenStudy (anonymous):

@Hero thanks so much

OpenStudy (lynfran):

here the elimination method using 1st and 2nd equation subtract we get y-z=32-39 now using the 3rd equation subtract y-z=-7 y+z=47 we get -z-z=-7-47 -2z=-54 z=-54/-2 z=27 now sub z=27 into 3rd equation we get y+z=47 y=47-27 y=20 now sub y=20 in 1st equation we get x+y=32 x+20=32 x=32-20 x=12

OpenStudy (anonymous):

this method seems confusing by the subtraction method seems easier to follow

hero (hero):

That WAS the subtraction method. Lyn just wrote it awkwardly

OpenStudy (anonymous):

ohh i c

OpenStudy (lynfran):

no thats the elimination method in details

hero (hero):

Subtraction/Elimination Method

hero (hero):

Interchangeable description

OpenStudy (anonymous):

ohh i understand both when i see the steps which lyn has shown nicely